OpenStudy (anonymous):

What value of n makes the equation true? n – 27.3 = 15.9 A. 11.4 B. 12.4 C. 42.2 D. 43.2

OpenStudy (anonymous):

d right

geerky42 (geerky42):

Try to modify the equation so n would be isolated. Try to retriceboth sides by 27.3. Like this: \(n-27.3\color{red}{+27.3}=15.9\color{red}{+27.3}\\n\cancel{-27.3}+\cancel{27.3} = 15.9+27.3\\n = 15.9+27.3\\n=\cdots~?\)

geerky42 (geerky42):

Yep D

geerky42 (geerky42):

add both sides by 27.3*